Hi, I'm having a lot of trouble with a couple of different questions. I did do the homework on this, but I feel like I did not do an example like this.
A) (25u^2)^5/2

B) b^2/3b^2/3 / -b^1/3
C) Find the perimeter of the trapezoid. This one they gave us a problem like it, but they never made us set it up fully. So I'm not even sure how to set it up.
Sides: 4 square root of 50, 2 square root of 72, 3 square root of 2, & the square root of 50.

Answer by stanbon(75887)

A) (25u^2)^(5/2)
-----
The 2 in the exponent denominator means "take the square root".
The 5 in the exponent numerator means "raise to the 5th power".
------
(25u^2)^(5/2) = (5u)^5 = 3125u^5
======================================
B) b^2/3b^(2/3) / -b^(1/3)
------------
= [b^(4/3)/3] / [-b^(1/3)
-----
= [b^(4/3)/[-3b^(1/3)]
----
= b/(-3)
---
= (-1/3)b
=============================

C) Find the perimeter of the trapezoid. This one they gave us a problem like it, but they never made us set it up fully. So I'm not even sure how to set it up.
Sides: 4 square root of 50, 2 square root of 72, 3 square root of 2, & the square root of 50.
----
P = 2sqrt(50) + 2sqrt(72)+3sqrt(2)+sqrt(50)
----
P = 2*5sqrt(2) + 2*6sqrt(2) + 3 sqrt(2) + 5sqrt(2)
-----
P = 30sqrt(2)
======================
Cheers,
Stan H.
=================